OpenStudy (anonymous):

The half-life of carbon-14 is 5,730 years. How many grams of parent material will remain from a 200-gram sample after 3 half-lives? 100 50 25 12.5

OpenStudy (anonymous):

one half life gives \(\frac{1}{2}\) of the original amount two half lives gives \(\frac{1}{4}\) of the original amount three half lives gives \(\frac{1}{8}\) of the original amount

OpenStudy (anonymous):

whats the answer though? 100 50 25 12.5